how many ounces of a 5% alcohol solution must be mixed with 14 ounces of a 10% alcohol solution to make a 6% alcohol solution?
Answer by mananth(16946)

----------- percent ---------------- quantity
Alcohol I 10 ---------------- 14 ounces
AlcoholII 4 ---------------- x ounces
Total 6 ---------------- 14 + x ounces

10*14+4* =6(14+x)

140+4x=84+6x
4x-6x=84 -140

-2 x = -56

/ -2

x = 28 ounces 4 % Alcohol II